OpenStudy (anonymous):

Find the Radius of Convergence and the Interval of Convergence for the series. SUM x^n/(n3^n) as n=1 to inf

OpenStudy (anonymous):

\[\sum_{1}^{\inf} x ^{n}/(n3^{^{n}})\]

OpenStudy (anonymous):

ratio test gives me x/3n+3

OpenStudy (anonymous):

now from there is where i get lost...as I figure that limit is 0 so I can't figure the radius of convergence or interval.

OpenStudy (anonymous):

then you win!

OpenStudy (anonymous):

how would i express that though like what would i say ? this is the first time i had a problem like this.

OpenStudy (anonymous):

if limit is zero for all \(x\) then radius of convergence is \((-\infty, \infty)\) i.e. it converges for all \(x\)

OpenStudy (anonymous):

if the limit was say 3 then the radius of convergence would be \(\frac{1}{3}\)

OpenStudy (anonymous):

would the interval of convergence be the same...as I imagine you can't test -inf and inf. so R = AllReal and Interval = ALlLREAL ?

OpenStudy (anonymous):

all real yes

OpenStudy (anonymous):

but i am not sure your ratio is right. i mean your answer is right, but your ratio may be off

OpenStudy (anonymous):

I'll double check it... maybe I messed up some algebra.

OpenStudy (anonymous):

\[x ^{n+1}/((n+1)3^{n+1}) * n3^{n}/x ^{n}\]

OpenStudy (anonymous):

who scratch that!! the radius of convergence is not \((-\infty, \infty)\) !!!

OpenStudy (anonymous):

by was i ever wrong, sorry

OpenStudy (anonymous):

oh crap lol...if your wrong then im in big trouble lol.

OpenStudy (anonymous):

\[\frac{n3^nx^{n+1}}{(n+1)3^nx^n}=\frac{nx}{3(n+1)}\]

OpenStudy (anonymous):

\[\lim_{n\to \infty}\frac{n}{3(n+1)}=\frac{1}{3}\] radius of convergence is 3

OpenStudy (anonymous):

Doh! it was my fault, assuming I had the ratio right. I can take it from there. Now I also know what to do in the event if I do have something that = 0.

OpenStudy (anonymous):

Thanks satellite73. Your always helping me out and explaining lol :-). I may very well pass this calc course because of you.

OpenStudy (anonymous):